# Advancements in Visual Language Models for Remote Sensing: Datasets, Capabilities, and Enhancement Techniques

### v2

@@ 1 Introduction
Remote sensing technology can help people obtain various observation data from a distance, and it is widely applied in fields such as disaster monitoring [1], urban planning [2], and agricultural management [3], surface observation [4], environmental monitoring [5], etc. Consequently, remote sensing image processing technology has been a key topic of focus for researchers both domestically and internationally. Over the past decade, with the development of artificial intelligence technology, significant breakthroughs have also been made in remote sensing image processing techniques. Specifically, various 3D convolutional neural networks (CNNs) [6] have been specifically designed for hyperspectral and multispectral image classification, as well as spectral reconstruction. U-Net [7] is utilized for semantic segmentation and land cover mapping in aerial imagery. A pyramid network has been proposed for multi-scale object detection in SAR images. The YOLO framework [8] is applied for small target detection in infrared remote sensing images. Besides, AI techniques are also applied in fields such as denoising, fusion, enhancement, and compression of remote sensing images, achieving remarkable results. In summary, AI techniques have significantly advanced the processing of remote sensing images, elevating the capabilities in these fields. However, most of these methods primarily focus on image processing, and the models they employ are discriminative models. There are inherent limitations to this type of technology, such as the inability to incorporate some human common sense, and the trained models can only perform a single vision task.
@@
Recently, due to advancements in data availability, computational power, and the introduction of the transformer model [9], natural language processing (NLP) has made significant breakthroughs. By feeding tens of billions of text tokens and scaling the model to tens of billions of parameters, ChatGPT [10] achieved amazing performance in natural language understanding, language translation, text generation, question answering, etc. The impressive success of ChatGPT [10] has reignited enthusiasm for AI. Larger and more advanced models continue to emerge, showcasing unprecedented advancements in language understanding through their extensive knowledge and sophisticated reasoning abilities, demonstrating human-like thinking. Specifically, Google released the encoder-only model BERT [11], which pushed the GLUE [12] benchmark to 80.4% and achieved an accuracy of 86.7% on MultiNLI [13] and set a new record of 93.2 for the F1 score on the SQuAD v1.1 [14] question-answering test, surpassing human performance by 2.0 points. Meta released LLaMA [15], which contains a series of models for 7 billion parameters to 65 billion parameters and it can beat GPT-3 [16] with 1/10 parameters. The development of large language models reveals new possibilities for the advancement of artificial intelligence technology, particularly in generative AI. Compared to discriminative AI, generative AI models approach problem-solving in a more flexible and open manner, allowing for the modeling of a broader range of issues. The development of LLMs also offers new insights for addressing the challenges faced by discriminative models in the computer vision field.
@@
The introduction of models such as CLIP [17] has effectively bridged the modality gap between images and language, creating opportunities for innovative integration. CLIP [17], with its innovative multimodal encoder-decoder structure utilizing BERT [11] as a text encoder and Vision Transformer (ViT) [18] / ResNet [19] as an image encoder, exemplifies the synergy between images and language vocabulary, leading to improved accuracy in classification and detection tasks through a deep understanding of the relationship between image content and textual vocabulary . GLIP [20] further enhances this consistency between language sentences and image content. Furthermore, the emergence of VLM has fundamentally addressed numerous constraints associated with traditional discriminative models in remote sensing image processing, thereby catalyzing the transition to AI 2.0. Generally speaking, VLM is a type of model designed to integrate and understand both visual and textual information, enabling tasks such as image captioning, visual question answering, and cross-modal retrieval by leveraging the relationships between images and language. Compared to previous discriminative models, VLM offers greater flexibility in modeling various AI tasks, allowing it to handle multiple tasks within a single framework. Additionally, its multimodal perceptual ability brings it closer to human perception. In the past two years, both LLMs and VLMs have developed rapidly, with various stronger models being introduced, such as LLaMA [15], LLaVA [21], and GPT-4 [22], among others. The performance of related tasks has also been continually improved.
@@
Similar to previous discriminative AI technologies, which have significantly improved the field of remote sensing data processing, VLM—one of the core technologies in AI 2.0—also brings new opportunities for this domain. VLMs, by leveraging their multimodal capabilities, have demonstrated significant progress across various tasks within remote sensing, including geophysical classification [23, 24], object detection [25, 26], and scene understanding [27, 28]. By introducing LLaVA-1.5 [21] into the remote sensing field, Kuckreja et al. proposed GeoChat [25], which not only answers image-level queries but also accepts region inputs for region-specific dialogue [25]. Deng et al. designed ChangeChat, which utilizes multimodal instruction tuning to handle complex queries such as change captioning, category-specific quantification, and change localization [29]. Inspired by GPT-4 [22], Hu et al. developed a high-quality Remote Sensing Image Captioning dataset (RSICap) to facilitate the advancement of large VLMs in the remote sensing domain [30]. However, overall, the application of VLM technology in the field of remote sensing is still in its early stages. Current VLM-based remote sensing data processing technologies have already demonstrated significant potential, but there are still many issues that need to be addressed.
@@
Contribution. In this paper, we present a comprehensive review of VLM-based remote sensing applications, offering a thorough delineation of the principal methodologies and the most recent advancements in VLMs, particularly within the context of remote sensing. By providing an in-depth introduction to the most representative and influential works in this domain, the survey enables researchers to quickly acquire a nuanced understanding of the developmental trajectory of these methodologies, facilitating the classification and evaluation of different approaches. Furthermore, the survey extends its contribution by presenting a detailed classification of enhancement techniques, focusing on key components such as vision encoders, text encoders, and the alignment between vision and language. This classification not only systematically illustrates the research progress made within this specialized field but also serves as a valuable reference point, supporting and inspiring further innovative research endeavors by scholars aiming to explore and advance the boundaries of VLM applications in remote sensing.
@@
Survey Pipeline This survey is organised as follows. The section 2 presents an convice introduction to foundation models, including Transformer, Vision Transformer and the famious VLM—LLaVA. The section 3 outline the datasets used in VLM for RS, which can be further divided manually into collected data, data combining with existing dataset and data enhanced by VLMs and LLMs. The section 4 describe the capabilities of VLMs, specifically introduce the pure visual tasks and visual language tasks. In Section 5, current VLM-based remote sensing data processing methods are introduced according to three major improvement directions, with a comparison and analysis of different approaches. Finally, several promising future directions are discussed in Section 6.
Figure 1: The illustration of Transformer [9]
@@ 2.1 Transformer
where $d_{k}$ is a scaling factor, typically set to the dimension of the input feature head. Q, K, and V are variations of the input data, usually generated by a linear layer. For instance, $Q=W_{q}(x)$, $K=W_{k}(x)$, $V=W_{v}(x)$, where $x$ donates the input feature and $W_{q}$$W_{k}$ and $W_{v}$ are learnable parameters of the linear layers. When Q,K and V are variations of the same input feature, it is referred to as self-attention. If Q varies from $x_{1}$ while K and V vary from $x_{2}$ (with $x_{1}$ and $x_{2}$ coming from different sources), it is called cross-attention.
@@ 2.2 Vision Transformer
Inspired by the success and popular of transformer used in NLP, the Vision Transformer (ViT) model was proposed later and tailored for computer vision field. As shown in Figure 2, ViT is composed of three components, that is Linear Projection of Flattened Patches(Embedding layer), Transformer Encoder and MLP Head. In summary, to enable NLP models to process images, the images are serialized by cropping them into uniform-sized patches. These patches are then transformed into features using a projection layer. To ensure the model is sensitive to the positional information of the image patches, position embeddings are added to the corresponding patch embeddings. Currently, there are four types of position embedding methods, including 1) absolute position embedding; 2) relative position embedding (RPE) [31]; 3) conditional position embedding (CPE) [32]; 4) rotary position embedding (RoPE) [33]. The ViT adopts the first one.
Compared to convolutional networks, ViT has a significant advantage in scaling up ability. Its performance on image processing does not saturate quickly as the model size increases, allowing for better utilization of larger models and enhanced capabilities in handling more complex tasks. This characteristic has sparked a surge of research interest in the ViT architecture, leading to the continuous emergence of various improved versions of ViT models, such as SwinTransformer [34], Deit [35], etc. So far, the model paradigms in both the NLP and CV fields have been unified, which has laid the foundation for the emergence of VLM.
@@ 2.3 Vision Language Model
The introduction of the Transformer has significantly advanced language processing, while the emergence of ViT has unified the processing paradigms in NLP and CV. CLIP [17] further broke down the modal barriers between language and images. Building on these foundations, VLMs have been proposed by merging ViT and LLMs. Figure 3 shows the architecture of LLaVA [21], a very representive VLM model. From Figure 3, we can see that the core idea of LLaVA is to combine visual tokens with text tokens and feed them into a large language model (LLM) to generate text-based answers. Pre-trained CLIP visual encoder is adopted here to extract visual tokens, and a trainable projection layer is build to convert visual tokens to language embedding tokens.
@@ 3 Datasets
Figure 5: Comparison between traditional datasets and VLM dataset in remote sensing.
@@
  • Combining datasets, which are build by combing several existing datasets and adding part new language annotations.
@@
In this section, we will delve into the construction methods and applicable task domains of existing remote sensing VLMs datasets, offering insights into how these datasets support extensive applications and in-depth research within the field. The statistical infromation of current existing datasets are summarized in Table 1
Method Dataset Source Image Image Size Task
Manual HallusionBench [37] - 346 - VQA
RSICap [30] DOTA 2585 512 IC
CRSVQA [38] AID 4639 600 VQA
Combining with Existing Datasets SATIN [39] Million-AID,WHU-RS19, SAT-4,AID1 $\approx$775K - SC
GeoPile [40] NAIP,RSD46-WHU, MLRSNet,RESISC45, PatternNet 600K - -
SatlasPretrain [41] UCM,BigEarthNet, AID,Million-AID, RESISC45,FMoW, DOTA,iSAID 856K 512 -
RSVGD [42] DIOR 17402 800 VG
RefsegRS [43] SkyScapes 4420 512 RRSIS
SkyEye-968K [44] RSICD,RSITMD, RSIVQA,RSVG1 968K - -
MMRS-1M [45] AID,RSIVQA, Syndney-Captions1 1M - -
RSSA [46] DOTA-v2,FAIR1M 44K 512 -
FineGrip [47] MAR20 2649 - -
RRSIS-D [48] - 17402 800 RRSIS
RingMo [49] Gaofen,GeoEye, WorldView,QuickBird1 2096640 448 -
GRAFT [50] NAIP,Sentinel-2 - - -
SkySense [51] WorldView,Sentinel-1, Sentinel-2 21.5M - -
EarthVQA [52] LoveDA,WorldView 6000 - VQA
GeoSense [53] Sentinel-2,Gaofen, Landsat,QuickBird $\approx$9M 224 -
Enhenced by VLMs and LLMs RS5M [26] LAION2B-en, LAION400M, LAIONCOCO1 5M - -
SkyScript [27] Google Earth Engine, OpenStreetMap 2.6M - -
LHRS-Align [54] Google Earth Engine, OpenStreetMap 1.15M - -
GeoChat [25] SAMRS,NWPU-RESISC-45, LRBEN,Floodnet 318k - -
GeoReasoner [55] Google Street View, OpenStreetMap 70K+ - StreetView
HqDC-1.4M [46] Million-AID, CrowdAI, fMoW,CVUSA, CVACT,LoveDA $\approx$1.4M 512 -
ChatEarthNet [56] Sentinel-2,WorldCover 163488 256 -
VRSBench [57] DOTA-v2,DIOR 29614 512 -
FIT-RS [58] STAR 1800.8K 512 -
1 Only a subset of common datasets are enumerated.
@@ 3.1 Manual Datasets
HallusionBench [37], the first benchmark designed to examine visual illusion and knowledge hallucination of large visual language models(LVLMs) and analyze the potential failure modes based on each hand-crafted example pair, consists of 455 visual-question control pairs, including 346 different figures and a total of 1129 questions on diverse topics. RSICap [30], RSIEval comprises five experts annotated captions with rich and high-quality information, thus the model achieves satisfactory performance with fine-tuning. CRSVQA [38] employs annotators from different backgrounds, including geography, surveying, mapping, to write complex questions that should reason the relationship between the objects or a concise question related to the image content. This can effectively prevent the problems of data redundancy and language bias in existing RSVQA datasets [59].

### v3

@@ 1 Introduction
Remote sensing technology can help people obtain various observation data from a distance, and it is widely applied in fields such as disaster monitoring [1], urban planning [2], agricultural management [3], surface observation [4], environmental monitoring [5], etc. Consequently, remote sensing image processing technology has been a key topic of focus for researchers both domestically and internationally. Over the past decade, with the development of artificial intelligence technology, significant breakthroughs have also been made in remote sensing image processing techniques. Specifically, various 3D convolutional neural networks (CNNs) [6] have been specifically designed for hyperspectral and multispectral image classification, as well as spectral reconstruction. U-Net [7] is utilized for semantic segmentation and land cover mapping in aerial imagery. A pyramid network has been proposed for multi-scale object detection in Special Administrative Region (SAR) images. The YOLO framework [8] is applied for small target detection in infrared remote sensing images. Additionally, AI techniques are also applied in fields such as denoising, fusion, enhancement, and compression of remote sensing images, achieving remarkable results. For specific arts in remote sensing, FFCA-YOLO [9] designed three plug-and-play modules, which are used to enhance and fuse features in the network, as well as to be spatial context aware, in order to improve local region perception and global correlation as much as possible without increasing complexity. Zhang et al. [10] improved YOLOv5’s perception and detection by integrating spatial-to-depth (SPD) elements, and the CoTC3 module to utilize contextual information. Liu et al. [11] proposed the MLPA method, which enhances cross-domain hyperspectral imaging (HSI) classification by aligning multi-level features to improve generalization and discriminative ability in the target domain. LBA-MCNet [12] enhanced ORSI-SOD by integrating localization, balancing, and affinity to improve boundary detection and foreground–background context modeling. MSC-GAN [13] leveraged MSC architecture, GFR modules, and channel enhancement to improve multitemporal cloud removal through efficient feature interaction and fusion. In summary, AI techniques have significantly advanced the processing of remote sensing images, elevating the capabilities in these fields. However, most of these methods primarily focus on image processing, and the models they employ are discriminative models. There are inherent limitations to this type of technology, such as the inability to incorporate some human common sense, and the trained models can only perform a single vision task.
@@
Recently, due to advancements in data availability, computational power, and the introduction of the Transformer model [14], natural language processing (NLP) has made significant breakthroughs. By feeding tens of billions of text tokens and scaling the model to tens of billions of parameters, ChatGPT [15] achieved amazing performance in natural language understanding, language translation, text generation, question answering, etc. The impressive success of ChatGPT [15] has reignited enthusiasm for AI. Larger and more advanced models continue to emerge, showcasing unprecedented advancements in language understanding through their extensive knowledge and sophisticated reasoning abilities, demonstrating human-like thinking. Specifically, Google released the encoder-only model BERT [16], which pushed the GLUE [17] benchmark to 80.4% and achieved an accuracy of 86.7% on MultiNLI [18] and set a new record of 93.2 for the F1 score on the SQuAD v1.1 [19] question-answering test, surpassing human performance by 2.0 points. Meta released LLaMA [20], which contains a series of models for 7 billion parameters to 65 billion parameters, and it can beat GPT-3 [21] with 1/10 parameters. The development of large language models reveals new possibilities for the advancement of artificial intelligence technology, particularly in generative AI. Compared to discriminative AI, generative AI models approach problem-solving in a more flexible and open manner, allowing for the modeling of a broader range of issues. The development of large language models (LLMs) also offers new insights for addressing the challenges faced by discriminative models in the computer vision field.
@@
The introduction of models such as CLIP [22] has effectively bridged the modality gap between images and language, creating opportunities for innovative integration. CLIP [22], with its innovative multimodal encoder–decoder structure utilizing BERT [16] as a text encoder and Vision Transformer (ViT) [23]/ResNet [24] as an image encoder, exemplifies the synergy between images and language vocabulary, leading to improved accuracy in classification and detection tasks through a deep understanding of the relationship between image content and textual vocabulary. GLIP [25] further enhances this consistency between language sentences and image content. Furthermore, the emergence of VLM has fundamentally addressed numerous constraints associated with traditional discriminative models in remote sensing image processing, thus catalyzing the transition to AI 2.0. Generally speaking, VLM is a type of model designed to integrate and understand both visual and textual information, enabling tasks such as Image Captioning, Visual Question Answering, and crossmodal retrieval by leveraging the relationships between images and language. Compared to previous discriminative models, VLM offers greater flexibility in modeling various AI tasks, allowing it to handle multiple tasks within a single framework. Additionally, its multimodal perceptual ability brings it closer to human perception. In the past two years, both LLMs and VLMs have developed rapidly, with various stronger models being introduced, such as LLaMA [20], LLaVA [26], and GPT-4 [27], among others. The performance of related tasks has also been continually improved.
@@
Similarly to previous discriminative AI technologies, which have significantly improved the field of remote sensing data processing, VLMs—one of the core technologies in AI 2.0—also bring new opportunities for this domain. VLMs, by leveraging their multimodal capabilities, have demonstrated significant progress across various tasks within remote sensing, including geophysical classification [28, 29], object detection [30, 31], and scene understanding [32, 33]. By introducing LLaVA-1.5 [26] into the remote sensing field, Kuckreja et al. proposed GeoChat [30], which not only answers image-level queries but also accepts region inputs for region-specific dialogue [30]. Deng et al. designed ChangeChat, which utilizes multimodal instruction tuning to handle complex queries such as change captioning, category-specific quantification, and change localization [34]. Inspired by GPT-4 [27], Hu et al. developed a high-quality Remote Sensing Image Captioning dataset (RSICap) to facilitate the advancement of large VLMs in the remote sensing domain [35]. Wang et al. [36] proposed a zero-shot target recognition framework for SAR, which combines a vision–language model with a diffusion generative model to achieve the construction of 3D models. Changen2 [37] is a generative model that creates time-series images with semantic and changeable labels from single-temporal images for collecting and annotating large multitemporal remote sensing datasets.
The publication growth trend in Figure 1 illustrates the development timeline of Visual–Language Models (VLMs) in the remote sensing (RS) field. Growth began after the release of CLIP [22], and in early 2023, following the successive releases of LLaVA [26] and GPT-4 [27], numerous high-performing VLMs started to emerge in the RS domain.
Figure 1: Number of publications for VLMs in RS (from Web of Science).
In practical applications, VLMs are beginning to demonstrate their potential in the remote sensing domain, albeit still in the early stages of adoption. A prominent example is GeoForge [38], which utilizes a customized GPT-4 model trained with additional geospatial vocabulary and knowledge to handle tasks like mapping administrative boundaries and geocoding major cities. For instance, users can interact with GeoForge to easily obtain the administrative boundaries of Bangladesh and the names of its five most important cities, and can further generate geocoded markers for these five cities based on the obtained information. Despite these advancements, the range of practical applications remains limited, and widespread industry implementation is yet to be realized. This early stage underscores the significant potential of VLMs like GeoForge to enhance remote sensing workflows while highlighting the need for ongoing development to overcome existing challenges and expand their real-world applicability.
Beyond these direct applications, VLMs also demonstrate potential advantages over traditional models in handling long-tail distributions and supporting multi-task learning. Pre-trained on extensive and diverse datasets, VLMs can achieve strong generalization capabilities, enabling them to effectively generalize to rare categories and thus mitigate the challenges associated with long-tail distributions. In addition, thanks to their generative output paradigm, VLMs can simultaneously tackle multiple tasks—such as classification, question answering, and counting—within a single framework.
@@
In this paper, we present a comprehensive review of VLM-based remote sensing applications; to ensure coverage of relevant research, we utilized Google Scholar and employed keywords such as “Visual language models for remote sensing”, “Multimodal remote sensing datasets”, and “Deep learning for remote sensing imagery”, offering a thorough delineation of the principal methodologies and the most recent advancements in VLMs, particularly within the context of remote sensing. By providing an in-depth introduction to the most representative and influential works in this domain, the survey enables researchers to quickly acquire a nuanced understanding of the development trajectory of these methodologies, facilitating the classification and evaluation of different approaches. Furthermore, the survey extends its contribution by presenting a detailed classification of enhancement techniques, focusing on key components such as vision encoders, text encoders, and alignment between vision and language. This classification, which differs from existing surveys categorized by task type [39], not only systematically illustrates the research progress made within this specialized field but also serves as a valuable reference point, supporting and inspiring further innovative research endeavors by scholars aiming to explore and advance the boundaries of VLM applications in remote sensing.
@@
This survey is organized as shown in Figure 2. Section 2 presents a convincing introduction to foundation models, including Transformer, Vision Transformer, and the famous VLM—LLaVA. Section 3 outlines the datasets used in VLM for RS, which can be further divided manually into collected data, data combined with existing datasets, and data enhanced by VLMs and LLMs. Section 4 describes the capabilities of VLMs, specifically introducing the pure visual tasks and vision–language tasks. In Section 5, current VLM-based remote sensing data processing methods are introduced according to three major improvement directions, with a comparison and analysis of different approaches. Finally, several promising future directions are discussed in Section 6.
Figure 2: Organization of this survey.
Figure 3: An illustration of Transformer [14].
@@ 2.1 Transformer
where $d_{k}$ is a scaling factor, typically set to the dimension of the input feature head. $Q$, $K$, and $V$ are variations of the input data, usually generated by a linear layer. For instance, $Q=W_{q}(x)$, $K=W_{k}(x)$, $V=W_{v}(x)$, where $x$ denotes the input feature and $W_{q}$, $W_{k}$ and $W_{v}$ are learnable parameters of the linear layers. When Q, K and V are variations of the same input feature, it is referred to as self-attention. If Q varies from $x_{1}$ while K and V vary from $x_{2}$ (with $x_{1}$ and $x_{2}$ coming from different sources), it is called cross-attention.
@@ 2.2 Vision Transformer
Inspired by the success and popularity of transformer used in NLP, the Vision Transformer (ViT) model was proposed later and tailored for the computer vision field. As shown in Figure 4, ViT is composed of three components: Linear Projection of Flattened Patches (embedding layer), Transformer Encoder, and MLP Head. In summary, to enable NLP models to process images, the images are serialized by cropping them into uniform-sized patches. These patches are then transformed into features using a projection layer. To ensure the model is sensitive to the positional information of the image patches, position embeddings are added to the corresponding patch embeddings. Currently, there are four types of position, embedding methods, including (1) absolute position embedding; (2) relative position embedding (RPE) [40]; (3) conditional position embedding (CPE) [41]; (4) rotary position embedding (RoPE) [42]. The ViT adopts the first one.
Compared to convolutional networks, ViT has a significant advantage in scaling-up ability. Its performance on image processing does not saturate quickly as the model size increases, allowing for better utilization of larger models and enhanced capabilities in handling more complex tasks. This characteristic has sparked a surge of research interest in the ViT architecture, leading to the continuous emergence of various improved versions of ViT models, such as SwinTransformer [43], Deit [44], etc. So far, the model paradigms in both the NLP and CV fields have been unified, which has laid the foundation for the emergence of VLMs.
@@ 2.3 Vision Language Model
The introduction of Transformers has significantly advanced language processing, while the emergence of ViT has unified processing paradigms in both NLP and CV. Subsequently, CLIP [22] breaks down the remaining barriers between language and images. Building on these foundations, VLMs have been proposed by combining ViT and LLMs. A representative paradigm among these VLMs is LLaVA [26], whose architecture is illustrated in Figure 5.
The core idea of LLaVA is to integrate visual tokens with textual tokens and feed them into an LLM to produce text-based answers. Specifically, the visual encoder extracts image features, and a projection layer converts these features into language-embedding tokens, bridging the gap between different modalities. These visual tokens are subsequently input into the LLM together with the text tokens, enabling multimodal interaction.
In practice, a pre-trained CLIP model is often employed as the visual encoder to extract visual tokens, because it has been trained extensively on large-scale image–text pairs. This training allows the CLIP visual encoder to map image features into a representation space comparable to that of text features, laying the groundwork for subsequent conversion into language-embedding tokens.
@@ 3 Datasets
Figure 7: Comparison of traditional (left) and VLM datasets (right) in remote sensing.
@@
  • Combined datasets, which are built by combining several existing datasets and adding part-new language annotations.
@@
In this section, we will delve into the construction methods and applicable task domains of existing remote sensing VLM datasets, offering insights into how these datasets support extensive applications and in-depth research within the field. The statistical information of current existing datasets are summarized in Table 1.
@@ 3.1 Manual Datasets
HallusionBench [46], the first benchmark designed to examine visual illusion and knowledge hallucination of large vision–language models (LVLMs) and analyze the potential failure modes based on each hand-crafted example pair, consists of 455 visual–question control pairs, including 346 different figures and a total of 1129 questions on diverse topics. RSICap’s [35] RSIEval comprises five expert-annotated captions with rich and high-quality information; thus, the model achieves satisfactory performance with fine-tuning. CRSVQA [47] employs annotators from different backgrounds, including geography, surveying, and mapping, to write complex questions that should reason the relationship between the objects or a concise question related to the image content. This can effectively prevent the problems of data redundancy and language bias in existing RSVQA datasets [48].
Method Dataset Year Source Image Image Size Task
Manual Datasets HallusionBench [46] 2023 - 346 - VQA
RSICap [35] 2023 DOTA 2585 512 IC
CRSVQA [47] 2023 AID 4639 600 VQA
Combining Datasets SATIN [49] 2023 Million-AID, WHU-RS19, SAT-4, AID 1 $\approx$775 K - SC
GeoPile [50] 2023 NAIP, RSD46-WHU, MLRSNet, RESISC45, PatternNet 600 K - -
SatlasPretrain [51] 2023 UCM, BigEarthNet, AID, Million-AID, RESISC45, FMoW, DOTA, iSAID 856 K 512 -
RSVGD [52] 2023 DIOR 17,402 800 VG
RefsegRS [53] 2024 SkyScapes 4420 512 RRSIS
SkyEye-968 K [54] 2024 RSICD, RSITMD, RSIVQA, RSVG 1 968 K - -
MMRS-1M [55] 2024 AID, RSIVQA, Syndney-Captions 1 1 M - -
RSSA [56] 2024 DOTA-v2, FAIR1M 44 K 512 -
FineGrip [57] 2024 MAR20 2649 - -
RRSIS-D [58] 2024 - 17,402 800 RRSIS
RingMo [59] 2022 Gaofen, GeoEye, WorldView, QuickBird 1 $\approx$2 M 448 -
GRAFT [60] 2023 NAIP, Sentinel-2 - - -
SkySense [61] 2024 WorldView, Sentinel-1, Sentinel-2 21.5 M - -
EarthVQA [62] 2024 LoveDA, WorldView 6000 - VQA
GeoSense [63] 2024 Sentinel-2, Gaofen, Landsat, QuickBird $\approx$9 M 224 -
Automatically Annotated Datasets RS5M [31] 2024 LAION2B-en, LAION400M, LAIONCOCO 1 5M - -
SkyScript [32] 2024 Google Earth Engine, OpenStreetMap 2.6 M - -
LHRS-Align [64] 2024 Google Earth Engine, OpenStreetMap 1.15 M - -
GeoChat [30] 2024 SAMRS, NWPU-RESISC-45, LRBEN, Floodnet 318 K - -
GeoReasoner [65] 2024 Google Street View, OpenStreetMap 70 K+ - StreetView
HqDC-1.4 M [56] 2024 Million-AID, CrowdAI, fMoW, CVUSA, CVACT, LoveDA $\approx$1.4 M 512 -
ChatEarthNet [66] 2024 Sentinel-2, WorldCover 163,488 256 -
VRSBench [67] 2024 DOTA-v2, DIOR 29,614 512 -
FIT-RS [68] 2024 STAR 1800.8 K 512 -
1 Only a subset of common datasets are enumerated.
However, manual datasets are time-consuming and expensive to create, requiring domain experts to annotate data. Even with expert annotators, inconsistencies and errors may arise, affecting dataset quality. Additionally, the limited capacity for manual annotation results in smaller datasets, which can hinder model generalization. To ensure consistency, clear annotation guidelines and multiple annotators for cross-validation are essential. Regular audits by experts can maintain quality. Using semi-automated tools can improve efficiency, helping manage large-scale manual annotation projects.